A survey of athletes at a high school is conducted, and the following facts are discovered: 34% of the athletes are football players, 58% are basketball players, and 14% of the athletes play both football and basketball. An athlete is chosen at random from the high school: what is the probability that the athlete is either a football player or a basketball player?
Probability =
% (Please enter your answer as a percent)

Answer by ikleyn(52781)

.

(1)  Question "either F or B" literally means  "(F or B separately, EXCLUDING F and B simultaneously)".


     Under this understanding,  P(either F or B) = (P(F) - P(both)) + (P(B) - P(both)) = 

                                                 = (0.34 - 0.14)    + (0.58 - 0.14) = 0.64 = 64%.    ANSWER




(2)  Question "F or B"  is close, but DIFFERENT.  It literally means  "(F or B, including F and B simultaneously)".


      Under this understanding,  P(F or B) = P(F) + P(B) - P(both) = 

                                           = 0.34 + 0.58 -  0.14   = 0.78 = 78%.    ANSWER



So, even small differences may lead to different results.


Therefore, the reader should be very attentive and very confident 
in the meaning of the question to answer it in a right way.

Solved.